# Context-Based Meta Reinforcement Learning for Robust and Adaptable Peg-in-Hole Assembly Tasks

### v3

@@ V Experimental Results
    The proposed modifications and methods enhance the real-world adaptation performance of PEARL in PiH assembly tasks and are applicable to different peg shapes.
@@
    The proposed OOD adaptation procedure enables PEARL to successfully and safely perform PiH assembly tasks with high uncertainty in the hole position.
[…]
The incremental peg movement in Cartesian coordinates as output of the meta RL agent is converted into joint torque commands via MoveIt [29] and ROS (Robot Operating System) [30]. The policy, context encoder, and alternative context encoder are multi-layer perceptron (MLPs) with three hidden layers of 200 neurons each. The latent space has two dimensions for tasks with uncertainty in the hole position and five dimensions for tasks with additional uncertainty in the hole orientation.
@@ V-A Effect of the Proposed Context Data Modification
Figure 10: Success rates of our modified PEARL (MP) agent and the original PEARL (OP) agent in tasks with uncertainty in the hole position and orientation. As can be seen, our MP agent significantly outperforms the OP agent in all tasks, and our MP agent is more sample-efficient than SAC.
@@ V-A1 Training Efficiency
We tested all agents after each epoch in all tasks for 10 trials, each trial consists of 10 successive adaptation trajectories, and calculated the average return. Return is the summation of rewards across the trajectory. The average return of each agent’s random seeds across the training epochs is represented by solid lines in Fig. 8, while the shaded regions represent the standard deviation. The results indicate that our MP agent achieves higher performance than the OP agent and requires less number of epochs to achieve its best performance. This proves that the modified context data enhances the training efficiency.
The degradation of our MP agent performance, especially in test tasks in Fig. 8(b) and Fig. 8(c), occurs because as the agent’s performance improves, it samples more context data inside the hole, which is unique to each task (as detailed in Sec. IV-A2). The context encoder relies on this unique data to infer the task and gradually neglects the context data outside the hole, which is similar across multiple tasks as shown in Fig. 6(a), by assigning high variance to it.
We trained a third agent (MR) that uses $m$ as a reward, instead of the negative $L_{2}$ distance, in addition to using it in the modified context data. The results in Fig. 8 show that using $m$ as a reward degrades the agent’s performance, that is why our MP agent uses $m$ in the context data only and uses the negative $L_{2}$ distance as a reward to train the policy.
@@ V-A2 Adaptation Efficiency
(a) Gear peg gear hole (GPGH).
(b) Gear peg circular shaft (GPCS).
@@
Figure 13: Success rates of our modified PEARL agents using a force/torque sensor and the original PEARL (OP) agents in the real-world GPGH and GPCS PiH assembly tasks. As can be seen, our MP agents outperform the OP agent in all the tasks and peg shapes (see Fig. 11).
@@ V-A3 Robustness to Uncertainty in Hole Orientation
We adapted the agents to handle additional uncertainty in the hole orientation, i.e., ±$\pm 15^{\circ}$ around the vertical axis. To achieve this, we introduced an additional dimension representing orientation into the estimated and actual hole positions, the peg position, and the action. Furthermore, we modified the reward function and the motion towards the actual hole to account for this orientation uncertainty. The success rates of the OP and MP agents after training for 100 epochs are represented in Fig. 10. The OP agent fails to learn the task achieving only 10% success rate, while our MP agent achieves 100 % success rate in all tasks, except for 90% success rate in test tasks with trajectory-based adaptation. Similar to our previous results, our MP agent is more sample-efficient during adaptation than the SAC agent trained from scratch. These results prove that the modified context data enhances the agent’s ability to infer tasks and its robustness to more sources of uncertainty. The results show that even if the reward is known, the OP agent will not succeed, which proves the importance of our context data modification.
V-B Real-World Adaptation Performance
After meta training in simulation, we tested the OP, using the sparse reward trajectory-based adaptation procedure in [4], and our MP agents in real-world assembly tasks with peg shapes of different complexity: Cuboid Peg in Square Hole (CPSH) (see Fig. 2), Gear-shaped Peg in Gear-shaped Hole (GPGH) (see Fig. 11(a)), and hollow Gear Peg in Circular-shaped gear Shaft (GPCS) (see Fig. 11(b)). The actual hole positions of the tasks relative to the estimated position are represented in Fig. 8(d). We adapt our MP agent to use force/torque sensor data using the method proposed in Sec. IV-C using 16,000 real-world samples for each of the three configurations, which is less than the data collected from one task per epoch during meta training, which is 32,000 samples.
V-C Out-of-Distribution Adaptation
@@
We applied the adaptation procedure proposed in Sec. IV-D to OOD tasks whose actual hole offsets are in the range of 10, 20, 30, and 40 mm from the estimated hole position. The number of steps sampled per each posterior distribution $N$ is 100 transitions. The success rate of the agent is represented in Fig. 14(a). The results show that the agent is able to achieve 100% success in tasks with actual hole offsets which are 10 times greater than those used during meta training in 7,000 adaptation steps. Examples of the posterior distribution evolution in different tasks are shown in Fig. 14(b), which show how the proposed method gradually and safely explores the latent space till success.
@@ VI Conclusion
Our experimental evaluation in simulation as well as in real-world experiments demonstrated the superior performance of our modifications compared to the original PEARL algorithm regarding the training efficiency, adaptation performance, robustness, and generalization capabilities.

### v4

@@ V Experimental Results
    The proposed modifications and methods enhance the real-world adaptation performance of PEARL in PiH assembly tasks while being applicable to different peg shapes.
@@
    The proposed OOD adaptation procedure enables PEARL to successfully and safely perform PiH assembly tasks with a high uncertainty in the hole position.
Our simulation experiments are carried out using MuJoCo [28] with known actual hole positions, similar to [4]. Thus, the meta RL agents have access to the accurate values of the reward $r$ and the motion towards the actual hole $m$. For real-world experiments, we use the UR5e robotic arm with a built-in force/torque sensor and the D405 RealSense camera mounted on the end-effector to capture images before and after the action and we use OpenCV [29] to detect the hole and peg features in the captured 2D images. The starting position of the peg is 10 mm above the given estimated hole position and we assume an uncertainty of $\pm$$4\text{\,}\mathrm{m}\mathrm{m}$ in the actual hole positions, Similar to [4].
The incremental peg movement in Cartesian coordinates as output of the meta RL agent is converted into joint torque commands via MoveIt [30] and ROS (Robot Operating System) [31]. The policy, the context encoder, and the alternative context encoder are multi-layer perceptrons (MLPs) with three hidden layers of 200 neurons each. The latent space has two dimensions for tasks with uncertainty in the hole position and five dimensions for tasks with additional uncertainty in the hole orientation.
@@ V-A Effect of the Proposed Context Data Modification
Figure 10: Success rates of our modified PEARL (MP) agent and the original PEARL (OP) agent in tasks with uncertainty in the hole position and orientation. Our MP agent significantly outperforms the OP agent in all tasks, and is more sample-efficient than SAC.
@@ V-A1 Training Efficiency
We tested all agents after each epoch in all tasks for 10 trials, each trial consists of 10 successive adaptation trajectories, and calculated the average return. Return is the summation of rewards across the trajectory. The average return of each agent’s random seeds across the training epochs is represented by solid lines in Fig. 8. The shaded regions represent the standard deviation. The results indicate that our MP agent achieves higher performance than the OP agent and requires less number of epochs to achieve its best performance. This proves that the modified context data enhances the training efficiency.
The degradation of our MP agent performance, especially in test tasks in Fig. 8(b) and Fig. 8(c), occurs because the agent samples more context data inside the hole, which is unique to each task (as detailed in Sec. IV-A2), as its performance improves. The context encoder relies on this unique data to infer the task and gradually neglects the context data outside the hole, which is similar across multiple tasks as shown in Fig. 6(a), by assigning a high variance to it.
We trained a third agent (MR) that uses $m$ as the reward, instead of the negative $L_{2}$ distance, in addition to using it in the modified context data. The results in Fig. 8 show that using $m$ as a reward degrades the agent’s performance. That is why our MP agent uses $m$ in the context data only and uses the negative $L_{2}$ distance as a reward to train the policy.
@@ V-A2 Adaptation Efficiency
(a) Gear peg/ Gear hole (GPGH).
(b) Gear peg/ Circular shaft (GPCS).
@@
Figure 13: Success rates of our modified PEARL agents using a force/torque sensor and the original PEARL (OP) agents in real-world GPGH and GPCS PiH assembly tasks. Our MP agents outperform the OP agent in all tasks and peg shapes (see Fig. 11).
@@ V-A3 Robustness to Uncertainty in Hole Orientation
We adapted the agents to handle additional uncertainty in the hole orientation, i.e., ± $\pm\,15^{\circ}$ around the vertical axis. To achieve this, we introduced an additional dimension representing orientation into the estimated and actual hole positions, the peg position, and the action. Furthermore, we modified the reward function and the motion towards the actual hole to account for this orientation uncertainty. The success rates of the OP and MP agents after training for 100 epochs are represented in Fig. 10. The OP agent fails to learn the task achieving only 10 % success rate, while our MP agent achieves 100 % success rate in all tasks, except for 90 % success rate in test tasks with the trajectory-based adaptation. Similar to our previous results, our MP agent is more sample-efficient during adaptation than the SAC agent trained from scratch. These results prove that the modified context data enhances the agent’s ability to infer tasks and its robustness to more sources of uncertainty. The results show that even if the reward is known, the OP agent will not succeed, which proves the importance of our context data modification.
@@
V-B Real-World Adaptation Performance
After meta training in simulation, we tested the OP, using the sparse reward trajectory-based adaptation procedure in [4], and our MP agents in real-world PiH with peg shapes of different complexity: Cuboid Peg in Square Hole (CPSH) (see Fig. 2), Gear-shaped Peg in Gear-shaped Hole (GPGH) (see Fig. 11(a)), and hollow Gear Peg in Circular-shaped gear Shaft (GPCS) (see Fig. 11(b)). The actual hole positions of the tasks relative to the estimated position are represented in Fig. 8(d). We adapt our MP agent to use force/torque sensor data using the method proposed in Sec. IV-C using 16,000 real-world samples for each of the three configurations, which is less than the data collected from one task per epoch during meta training, which is 32,000 samples.
V-C Out-of-Distribution Adaptation
We applied the adaptation procedure proposed in Sec. IV-D to OOD tasks whose actual hole offsets are in the range of 10, 20, 30, and $40\text{\,}\mathrm{m}\mathrm{m}$ from the estimated hole position. The number of steps sampled per each posterior distribution $N$ is 100 transitions. The success rate of the agent is represented in Fig. 14(a). The results show that the agent is able to achieve 100 % success in tasks with actual hole offsets which are 10 times greater than those used during meta training in 7,000 adaptation steps. Examples of the posterior distribution evolution in different tasks are shown in Fig. 14(b), which show how the proposed method gradually and safely explores the latent space till success.
@@ VI Conclusion
Our experimental evaluation in simulation as well as in the real world demonstrates the superior performance of our modifications compared to the original PEARL algorithm regarding the training efficiency, adaptation performance, robustness, and generalization capabilities.